Determine the principal P that must be invested at rate r = 9%, compounded monthly, so that $500,000 will be available for retirement in t = 9 years. (Round your answer to the nearest cent.)

Answer by lwsshak3(11628)

Determine the principal P that must be invested at rate r = 9%, compounded monthly, so that $500,000 will be available for retirement in t = 9 years. (Round your answer to the nearest cent.)
***
compound Interest formula: A=P(1+r/n)^nt, P=initial investment, r=interest rate, n=number of compounding periods per year, A=amt after t-years.
For given problem:
A=500,000
P=?
r=9%
n=12
t=9
..
500000=P(1+.09/12)^12*9
500000=P(1+.0075)^108
500000=P(1.0075)^108
P=500000/(1.0075)^108
P=223102.32
principal P that must be invested=$223,102.32